Given a triangle ABC, if A is 11 degrees, B is 49 degrees, and b is 45 cm, what is the length of a (round to one decimal.)
-----------------------------------------------------
What is the measure in degrees of the angle A = 7π / 6 radians?

Answer by Alan3354(69443)

Given a triangle ABC, if A is 11 degrees, B is 49 degrees, and b is 45 cm, what is the length of a (round to one decimal.)
---------------
Use the Law of Sines:
a/sin(A) = b/sin(B)
a = 45*sin(11)/sin(49)
a = 11.4
=================================
What is the measure in degrees of the angle A = 7π / 6 radians?
Degs = (180/pi)*radians
(7pi/6)*(180/pi) = 210 degs